OpenStudy (anonymous):

Which set of integers is not a Pythagorean triple? A. 12, 35, 37 B. 14, 46, 48 C. 16, 63, 65 D. 20, 99, 101

OpenStudy (ikram002p):

well u need to try them all a Pythagorean triple means numbers satisfy phthagorous which is \(\Huge c^2=a^2+b^2\) such that C is the biggest numbet so try \(\color{green }{\Huge 35^2 + 12^2 =37^2} \large\color{blue}{\text{ Pythagorean triple} }\) \(\color{red }{\Huge 14^2 + 46^2 \neq 48^2}\large\color{orange}{\text{ not Pythagorean triple} }\) \(\color{green }{\Huge 16^2 + 63^2 = 65^2}\large\color{blue}{\text{ Pythagorean triple} }\) \(\color{green }{\Huge 20^2 + 99^2 =101^2}\large\color{blue}{\text{ Pythagorean triple} } \)

OpenStudy (animalover02):

The answer is b for shore.